Bennie is three times older then tommie in seven years Bennie will be twice as old how old are they now
Found 2 solutions by oscargut, Alan3354:
Answer by oscargut(2103)

Bennie is three times older then tommie in seven years Bennie will be twice as old how old are they now

Let x = Bennie`s age now
Let y = Tommie`s age now
Then:
x = 3y
x+7 = 2(y+7)
3y+7 = 2y+14
y = 7
x = 21
Answer: Bennie has 21 years old and tommie 7 years old
you can ask me more at: mthman@gmail.com
Thanks

Answer by Alan3354(69443)

Bennie is three times older then tommie in seven years Bennie will be twice as old how old are they now
-----------------
3 times older is not the same as 3 times as old.
3 times older is 4 times as old.
================
B = Bennie`s age now
T Tommie`s age now
Then:
B = 4T
B + 7 = 2*(T + 7)
Sub for B
4T + 7 = 2T + 14
2T = 7
T = 3.5
B = 14